OpenStudy (anonymous):

given : sin (2A + 3B) = 1/3 and cos (2A - 3B) = 2^sqrt2/3, the value of sin 4A =...?

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

i am confuse...

OpenStudy (anonymous):

@tblue1998 can we write sin(4A) as sin [ {2A +3B} + {2A -3B} ] now we can apply sin(X+Y)= sinX*cosY + cosX*sinY

OpenStudy (anonymous):

is all well? @tblue1998

OpenStudy (anonymous):

yeah, thanks

OpenStudy (anonymous):

@niksva but, how to get the value of cos(2A+3B) ?

OpenStudy (anonymous):

use \[\sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

oh, you really help me...